One subunit of Pike Sports Company had the following financial results last month:

Pike-Subunit X Actual Results Flexible Budget Flexible Budget Variance (F or U)

% Variance

(F or U)

Direct Materials

$28,800 $27,000
Direct Labor 13,400 14,200
Indirect Labor 26,100 23,000
Utilities 12,200 11,300
Depreciation 29,000 29,000
Repairs and Maintenance 4,800 5,700
Total $114,300 $110,200

Requirements:

1.

Complete the performance evaluation report for this subunit. Enter the variance percent as a percentage of the budgeted amount rounded to two decimal places.

2.

Based on the data​ presented, what type of responsibility center is this​ subunit?

3.

Which items should be investigated if part of​ management's decision criteria is to investigate all variances exceeding​ $2,500 or​ 10%?

4.

Should only unfavorable variances be​ investigated? Explain.
Answer #1

Solution:

Pike-Subunit X Actual Results Flexible Budget Flexible Budget Variance (F or U)

% Variance

(F or U)

Direct Materials

$28,800 $27,000 6.67% U
Direct Labor 13,400 14,200 5.63% F
Indirect Labor 26,100 23,000 13.48% U
Utilities 12,200 11,300 7.96% U
Depreciation 29,000 29,000 0% No Variance
Repairs and Maintenance 4,800 5,700 15.79% F
Total $114,300 $110,200 3.72% U

Calculation of Percentage based on Flexible Budget Data

Actual

Flexible

Flexible Budget

Plke - Subunit X

Results

Budget

Variance $

F or U

Variance (F or U)

Direct Materials

$28,800

$27,000

$1,800

U

(1800/27,000) = 6.67%

U

Direct Labor

13,400

14,200

$800

F

(800 / 14,200) = 5.63%

F

Indirect Labor

26,100

23,000

$3,100

U

(3,100/23,000) = 13.48%

U

Utilities

12,200

11,300

$900

U

(900 / 11,300) = 7.96%

U

Depreciation

29,000

29,000

$0

No variance

0.00%

No Variance

Repairs and Maintenance

4,800

5,700

$900

F

(900 / 5,700) = 15.79%

F

Total

$114,300

$110,200

$4,100

U

(4,100 / 110,200) = 3.72%

U

2- The subunit is a cost center .

3-The following items should be investigated:-

Indirect labor

Repairs and maintenance

4-No, should be investigated to make sure they are not hurting the business in the long run.